timothys age in 11 years will be twice what it was 9 years ago
Answer by ewatrrr(24785)

x Timothy's age Now
x + 11 = 2(x-9)
x + 11 = 2x - 18
29 = x, Timothy's age Now
And...checking
40 = 2(20)